at 12 noon, the teacher sent home 1/3 of the students plus 6. At 1pm, she sent home 1/4 of the remaining students.At 2pm, she sent home the last 21 students. How many students were originally in school that day?
Answer by mananth(16949)

Total x
12 PM 1/ 3 x+ 6 were sent away
Remainder= x - 1/3x- 6
Balance in class 2/ 3x-6
1PM she sent away 1/ 2*( 2/3x-6)
Balance in class 2/3x-6 -(0.33x +3)
1/ 3 x-3
2.00PM She sent away 21
2 PM 1/3 x -3 -21 = 0
1/ 3 x- 24 = 0
1/ 3 x= 24
x= 72